Unit 19 Homework Assignment  Physics141-Summer 2009

Unit 19 homework after session 1 (Wednesday)

• Read chapter 22 in the text book ”Understanding Physics”.

• Work textbook problems 22-5, 22-13, 22-19. Unit 19 homework after session 2 (Friday)

• Work textbook problems 22-43, 22-45. Unit 19 homework after session 3 (Monday)

• Read chapter 23 in the text book ”Understanding Physics”. (section 23-9 is optional)

• Work textbook problems 23-4, 23-7, 23-24.

All Unit 19 homework problems and the unit 19 activity guide are due on the first day of Unit 20